# Mellivora Capensis: A Backdoor-Free Training Framework on the Poisoned Dataset without Auxiliary Data

### v2

@@ 5.4 Ablation Studies
5.4.1 Performance with Relabeling and Relearning
In our MeCa, after identifying the backdoor samples and clean samples, we relabel the backdoor samples and merge them with clean samples to obtain a clean and complete dataset, which is used to fine-tune the clean model then. Accordingly, we conduct experiments on the ResNet18 and CIFAR10 datasets to show the effect of the relabeling and relearning mechanism. Table IV shows the ACC and ASR of our MeCa after using relabeling and relearning. Compared with Table II and III, we can know that the ACC further improves, and the ASR also decreases a little. Specifically, after employing relabeling and relearning, the average ACC and ASR changes from $84.29\%$ to $86.65\%$ and from $0.33\%$ to $0.03\%$ against the $8$ backdoor attacks, respectively. We infer that the model learns better due to the increase of clean samples. More interestingly, we find that employing relabeling and relearning has a positive effect on clean-label backdoor attacks (e.g., SIG). The ASR drops from $2.64\%$ to $0.22\%$ after equipping the relabeling and relearning mechanism. We deduce that the efficacy of backdoor task learning is closely linked to the initial weights of the model. Therefore, a pretrained model can hardly be backdoored by poisoned samples, which also indicates model service providers use pretrained models as the base models and fine-tune them with untrusted data to avoid poisoning attacks.
@@ 5.4.1 Performance with Relabeling and Relearning
5.4.2 Effectiveness with Different Partition Rates
@@
TABLE V: Impact of poisoning ratio on defense performance (%).
Poison Ratio→ 0.1 0.15 0.2
@@
5.4.3 Effectiveness with Different Poisoning Ratios
In real-world application scenarios, it is impractical for the defender to access the poisoning ratios of the training data. Therefore, we also demonstrate that our MeCa is suitable for multiple poison ratios. Here, we experiment it on CIFAR10 against $7$ backdoor attacks, including BadNet, Blend, WaNet, PhysicalBA, TaCT, AdaptivaBlend, and AdaptivaPatch with poisoning ratios up to $0.1$, $0.15$, $0.2$. Note that SIG is not considered here because it is a clean-label attack that can just poison one-class samples. Moreover, the number of one-class samples is a maximum of $10.00\%$ in CIFAR10. That is, the poisoning ratios of SIG on CIFAR10 do not exceed $10.00\%$. Table V shows the experimental results. From Table V, we can know that with the variation of poisoning ratios, the ASR of our method is always $0.00\%$ against the $7$ backdoor attacks. In particular, we can find that the ASR is $0.66\%$ when the poisoning ratio is zero. We suppose that the reason is the misclassification of the model on the clean samples. In addition, the experimental results show that our method also maintains a fine ACC against different backdoor attacks even without employing relabeling and relearning methods.
5.4.4 Effectiveness with Different Models
To demonstrate the generalizability of our method, we conduct experiments on ResNet34 and MobileNetV2 with the poisoning ratio of $0.05$. Table VI shows the experimental results on CIFAR10. From Table VI, we can find that our MeCa works similarly well on ResNet34 and MobileNetV2. The ASR is always $0.00\%$ in the two model architectures. Moreover, the ACC also maintains a satisfactory performance (from $88.60\%$ to $88.70\%$ in ResNet34, from $84.33\%$ to $81.89\%$ in MobileNetV2). The experimental results indicate that our MeCa has good generalizability between different models.
@@ 5.4.4 Effectiveness with Different Models
5.4.5 Impact on clean and poisoned dataset
In real-world applications, it is impractical for the defender to know whether a given training dataset contains some poisoned samples. Therefore, to demonstrate the impact of our MeCa on the clean and poisoned dataset, we conduct experiments on CIFAR10 with the poisoning ratio of 00 and $0.05$. The experimental results (Table VII) illustrate that for training on the poisoned dataset, our MeCa can significantly reduce the average ASR (from $74.45\%$ to $0.33\%$) with a slight loss of ACC (from $85.81\%$ to $84.29\%$). On the other hand, when trained on a clean dataset, our MeCa has a minor impact on ACC (from $87.86\%$ to $86.04\%$). Note that the ASR is $0.66\%$ when training on the clean dataset with our MeCa. We infer that this can be attributed to the misclassification of the model on the clean samples. The experimental results further demonstrate that MeCa is a practical solution for dealing with unknown datasets.
6 Conclusion
In this paper, we first explore the relationship between the backdooor and perturbation by our theoretical analysis and experimental verification. Based on the results of the investigation, we propose a novel backdoor defense method (MeCa) to identify poisoned samples and train a clean model on a poisoned dataset. The proposed MeCa partitions the poisoned samples and clean samples according to their robustness to adversarial perturbation. There is no requirement for any auxiliary clean dataset or knowledge about the poisoned dataset (e.g., poisoning ratios) in the MeCa. Extensive experimental results show the superior performance of MeCa in defending against $8$ state-of-the-art backdoor attacks. Compared with the $7$ advanced backdoor defense methods, our MeCa has a lower ASR while maintaining a satisfactory ACC on the main task. In addition, the experimental results also demonstrate that the proposed MeCa has a fine generalization ability in different poisoning ratios and various model architectures. The relevant experimental results indicate our method has prominent potential and vital practicality for real-world application scenarios as well.

### v3

@@ V-D Ablation Studies
V-D1 Performance with Relabeling and Relearning
In our MeCa, after identifying the backdoor samples and clean samples, we relabel the backdoor samples and merge them with clean samples to obtain a clean and complete dataset, which is used to fine-tune the clean model then. Accordingly, we conduct experiments on the ResNet18 and CIFAR10 datasets to show the effect of the relabeling and relearning mechanism. Table IV shows the ACC and ASR of our MeCa after using relabeling and relearning. Compared with Table II and III, we can know that the ACC further improves, and the ASR also decreases a little. Specifically, after employing relabeling and relearning, the average ACC and ASR changes from $84.29\%$ to $86.65\%$ and from $0.33\%$ to $0.03\%$ against the $8$ backdoor attacks, respectively. We infer that the model learns better due to the increase of clean samples. More interestingly, we find that employing relabeling and relearning has a positive effect on clean-label backdoor attacks (e.g., SIG). The ASR drops from $2.64\%$ to $0.22\%$ after equipping the relabeling and relearning mechanism. We deduce that the efficacy of backdoor task learning is closely linked to the initial weights of the model. Therefore, a pretrained model can hardly be backdoored by poisoned samples, which also indicates that model service providers use pretrained models as the base models and fine-tune them with untrusted data to avoid poisoning attacks.
@@ V-D1 Performance with Relabeling and Relearning
V-D2 Effectiveness with Different Partition Rates
@@
TABLE V: Impact of poison ratio on defense performance (%).
Poison Ratio→ 10% 15% 20%
@@
V-D3 Effectiveness with Different Poison Ratios
In real-world application scenarios, it is impractical for the defender to access the poison ratios of the training data. Therefore, we also demonstrate that our MeCa is suitable for multiple poison ratios. Here, we experiment it on CIFAR10 against $7$ backdoor attacks, including BadNet, Blend, WaNet, PhysicalBA, TaCT, AdaptivaBlend, and AdaptivaPatch with poison ratios up to $0.1$, $0.15$, $0.2$. Note that SIG is not considered here because it is a clean-label attack that can just poison one-class samples. Moreover, the number of one-class samples is a maximum of $10.00\%$ in CIFAR10. That is, the poison ratios of SIG on CIFAR10 do not exceed $10.00\%$. Table V shows the experimental results. From Table V, we can know that with the variation of poison ratios, the ASR of our method is always $0.00\%$ against the $7$ backdoor attacks. In particular, we can find that the ASR is $0.66\%$ when the poison ratio is zero. We suppose that the reason is the misclassification of the model on the clean samples. In addition, the experimental results show that our method also maintains a fine ACC against different backdoor attacks even without employing relabeling and relearning methods.
V-D4 Effectiveness with Different Models
To demonstrate the generalizability of our method, we conduct experiments on ResNet34 and MobileNetV2 with the poison ratio of $5\%$. Table VI shows the experimental results on CIFAR10. From Table VI, we can find that our MeCa works similarly well on ResNet34 and MobileNetV2. The ASR is always $0.00\%$ in the two model architectures. Moreover, the ACC also maintains a satisfactory performance (from $88.60\%$ to $88.70\%$ in ResNet34, from $84.33\%$ to $81.89\%$ in MobileNetV2). The experimental results indicate that our MeCa has good generalizability between different models.
@@ V-D4 Effectiveness with Different Models
V-D5 Impact on clean and poisoned dataset
In real-world applications, it is impractical for the defender to know whether a given training dataset contains some poisoned samples. Therefore, to demonstrate the impact of our MeCa on the clean and poisoned dataset, we conduct experiments on CIFAR10 with the poison ratio of 0 and $5\%$. The experimental results (Table VII) illustrate that for training on the poisoned dataset, our MeCa can significantly reduce the average ASR (from $74.45\%$ to $0.33\%$) with a slight loss of ACC (from $85.81\%$ to $84.29\%$). On the other hand, when trained on a clean dataset, our MeCa has a minor impact on ACC (from $87.86\%$ to $86.04\%$). Note that the ASR is $0.66\%$ when training on the clean dataset with our MeCa. We infer that this can be attributed to the misclassification of the model on the clean samples. The experimental results further demonstrate that MeCa is a practical solution for dealing with unknown datasets.
VI Conclusion
In this paper, we first explore the relationship between the backdooor and perturbation by our theoretical analysis and experimental verification. Based on the results of the investigation, we propose a novel backdoor defense method (MeCa) to identify poisoned samples and train a clean model on a poisoned dataset. The proposed MeCa partitions the poisoned samples and clean samples according to their robustness to adversarial perturbation. There is no requirement for any auxiliary clean dataset or knowledge about the poisoned dataset (e.g., poison ratios) in the MeCa. Extensive experimental results show the superior performance of MeCa in defending against $8$ state-of-the-art backdoor attacks. Compared with the $7$ advanced backdoor defense methods, our MeCa has a lower ASR while maintaining a satisfactory ACC on the main task. In addition, the experimental results also demonstrate that the proposed MeCa has a fine generalization ability in different poison ratios and various model architectures. The relevant experimental results indicate our method has prominent potential and vital practicality for real-world application scenarios as well.